The cost of a taxicab is 0.95 plus 1.25 per mile.
If the fare was 28.45, how
Long was the ride?
If a limo service charges $15 for a 12- mile ride to the airport, which one is a better deal, cab or limo?

Answer by rwm(914)

y=mx+b
28.45=1.25x+.95
x is the number of miles
x=22
t=1.25*12+.95=15.95
l=15
the limo is more fun and cheaper